# Collective Outlier Detection and Enumeration with Conformalized Closed Testing

### v2

@@ Abstract
This paper develops a flexible distribution-free method for collective outlier detection and enumeration, designed for situations in which the presence of outliers can be detected powerfully even though their precise identification may be challenging due to the sparsity, weakness, or elusiveness of their signals. This method builds upon recent developments in conformal inference and integrates classical ideas from other areas, including multiple testing, rank tests, and non-parametric large-sample asymptotics. The key innovation lies in developing a principled and effective approach for automatically choosing the most appropriate machine learning classifier and two-sample testing procedure for a given data set. The performance of our method is investigated through extensive empirical demonstrations, including an analysis of the LHCO high-energy particle collision data set.
@@ 1.1 Background and Motivation
While conformal inference is gaining momentum in both statistics and machine learning, its focus in the context of outlier detection has so far primarily been on individual-level outlier identification, where each data point is separately evaluated as a potential outlier. However, individual-level detection is not always feasible. In many applications, outliers may be too rare or weak to reach statistical significance (Donoho and Jin, 2015), or they may exhibit unremarkable behavior in isolation but reveal anomalous patterns when analyzed collectively (Feroze et al., 2021), such as showing under-dispersion relative to inliers.
@@ 1.2 Contributions and Outline
We present ACODE, an Automatic Conformal Outlier Detection and Enumeration method. ACODE leverages two-sample rank tests applied to univariate conformity scores (Kuchibhotla, 2020), which can be generated by any model trained to distinguish outliers from inliers. By employing split-conformal inference, ACODE converts these scores into statistical tests controlling the type-I error rate. In addition, by utilizing the closed testing principle (Marcus et al., 1976), ACODE can construct simultaneous lower confidence bounds for the number of outliers within any subset of the test sample (Goeman and Solari, 2011). While these high-level ideas are intuitive, the flexibility of the conformal inference framework introduces substantial complexity and a great deal of implementation freedom. This brings us to the two key methodological questions addressed in this paper.
@@
Figure 1: Preview of performance of ACODE on the LHCO data. Left: 90% lower confidence bound for the number of outliers. Right: power against the global null hypothesis of no outliers at the 10% level (horizontal line). The results are shown as a function of the true number of outliers in a test set of cardinality 10,000. ACODE utilizes a testing procedure that may be adaptively selected (red curve) or fixed (other solid curves). Dotted curve: number of individual discoveries or power obtained by applying the Benjamini-Hochberg procedure (BH) to conformal p-values, controlling the false discovery rate below 10%.
As detailed in Section 3.2, our analysis of the LHCO data aims to (1) detect outliers, or unusual collision events, within a test sample, and (2) establish a 90% lower confidence bound for the number of these outliers. Our method not only controls errors for both tasks but also provides insightful inferences. This illustrates ACODE’s adaptability, as it selects the most suitable testing procedures in a data-driven manner. Moreover, ACODE can leverage a diverse toolbox of classification algorithms, automatically selecting an effective model for each scenario, choosing among both one-class and binary classifiers. This feature is particularly useful in practice, as the performance of different classifiers can vary substantially across scenarios, a point further elaborated in Section 3.
Additionally, Figure 1 highlights the challenge of conducting an adaptive data-driven analysis without incurring selection bias. It compares ACODE’s performance with that of a naive “cherry-picking” approach that greedily tests various classification algorithms and testing procedures, selecting the one with the most appealing outcomes. Unsurprising, such a heuristic can lead to inflated type-I errors. In contrast, ACODE is protected against selection bias while achieving similarly high power across all scenarios tested.
@@ 1.3 Related Work
Recent advancements in large-scale inference have led to methods for signal detection and estimation of the proportion of non-null effects (Cai and Sun, 2017). Donoho and Jin (2004) proposed an elegant solution for signal detection using Tukey’s Higher Criticism statistic. Meinshausen and Rice (2006) extended the Higher Criticism method to provide a lower confidence bound for the proportion of non-null effects. Goeman et al. (2021) further advanced this approach by embedding Higher Criticism into a closed testing procedure (Marcus et al., 1976), providing simultaneous bounds for the false discovery proportion (FDP). However, the FDP bounds derived from Higher Criticism are valid only under the assumption of independent p-values, which does not hold in the conformal inference setting.
Since the works by Genovese and Wasserman (2006) and Goeman and Solari (2011), there has been significant growth in methods providing simultaneous FDP bounds (Goeman et al., 2019; Blanchard et al., 2020; Katsevich and Ramdas, 2020; Tian et al., 2023; Heller and Solari, 2023). These have been utilized across various domains, including neuroimaging (Rosenblatt et al., 2018; Goeman et al., 2023) and genomics (Ebrahimpoor et al., 2020), as well as in permutation-based (Hemerik et al., 2019; Andreella et al., 2023; Blain et al., 2022; Vesely et al., 2023) and knockoff-based (Li et al., 2024) approaches.
@@
In this paper, we integrate conformal inference with closed testing to construct simultaneous lower confidence bounds for the number of outliers in any subset of the test set. A central challenge is selecting an effective local testing procedure suited to the data at hand. Recent findings show that a closed testing procedure is admissible if and only if all its local tests are admissible (Goeman et al., 2021), underscoring the importance of designing a closed testing approach with admissible local tests.
For local tests relying on the combination of independent p-values, Birnbaum (1954) established that there does not exist a general uniformly most powerful p-value combination method for all alternatives. Further, any (reasonable) combining method is admissible, in the sense that there exists an alternative for which it is optimal (see Heard and Rubin-Delanchy, 2018, for specific examples concerning standard combination methods).
When a uniformly most powerful test is unavailable, a possible strategy is to maximize power locally around the null hypothesis. Lehmann (1953) identified a nonparametric alternative where the Wilcoxon-Mann-Whitney (WMW) test (Wilcoxon, 1945; Mann and Whitney, 1947) is the locally most powerful rank test. Inspired by this optimality property, we will utilize the WMW rank test as a local test in the closed testing framework, although we will also consider several other options. Additionally, in Section 2.3.2, we will explore a natural extension of the WMW rank test, inspired by Shiraishi (1985).
@@
A typical concern with closed testing is the computational cost, generally exponential. However, efficient polynomial-time shortcuts are often available (Goeman et al., 2019; Dobriban, 2020; Tian et al., 2023). In Section 2.4 and in Appendix A3 we, respectively, present the closed testing method and the computational shortcuts that make our method feasible and scalable. As we will see, different local tests often require different shortcuts.
@@ 2.1 Setup and Problem Statement
[…]
In the following, we will refer to the data points sampled from $P_{0}$ as inliers and those from any $P_{j}\neq P_{0}$ as outliers. For each $j\in[n]$, define the null hypothesis $H_{j}:P_{j}=P_{0}$. Let $I_{0}:=\{j\in[n]:P_{j}=P_{0}\}$ denote the subset of inliers in the test set. Similarly, let $I_{1}=[n]\setminus I_{0}$ denote the subset of outliers in the test set.
Prior work on conformal inference for outlier detection (Bates et al., 2023; Marandon et al., 2024) has primarily focused on testing individual-level hypotheses $H_{j}$ for each $j\in[n]$, with the goal of controlling the FDR. However, as discussed in Section 1, there are many scenarios where the power to reject these individual-level hypotheses $H_{j}$ is very low, highlighting the need for an alternative approach.
In this paper, we address two key tasks related to collective outlier detection and enumeration. The first task, outlier detection, involves testing the potentially easier-to-reject composite hypothesis $H_{S}:=\cap_{j\in S}H_{j}$, which posits that a subset $S\subseteq[n]$ of test points contains no outliers. The second task, outlier enumeration, focuses on estimating the number of outliers within a subset $S\subseteq[n]$, specifically by providing a lower confidence bound, without necessarily identifying them. For both tasks, we consider cases where $S$ is either fixed or data-driven. In the latter scenario, we will utilize the closed testing principle (Marcus et al., 1976) to obtain simultaneous inferences that remain valid for all possible subsets $S$. This approach gives practitioners flexible tools for interactive data analysis.
@@ 2.2 Computing Conformity Scores and Conformal p-values
As long as $(X_{1},\ldots,X_{m},(Y_{i})_{i\in I_{0}})$ are exchangeable, $p_{j}$ is a valid p-value for $H_{j}$, in the sense that $\mathbb{P}[p_{j}\leq\alpha\mid H_{j}]\leq\alpha$ for all $\alpha\in[0,1]$. Further, $p_{j}$ is not too conservative, in the sense that $\mathbb{P}[p_{j}\leq\alpha\mid H_{j}]$ is not far from $\alpha$, as long as $m$ is large and the conformity scores are almost surely distinct; the latter would be a mild assumption (since one can always add a small amount of independent noise to $\hat{s}$) but is not required.
@@ 2.3.1 Existing Approaches
Combination methods that do not assume independence fall into three main categories: those valid for any type of dependence, which tend to be conservative in practice (Vovk and Wang, 2020); those designed for specific types of dependence, such as the PRDS assumption (Chen et al., 2023); and those based on permutations (see Stoepker et al., 2024, for a recent example involving the Higher Criticism statistic).
Combination methods with universal validity are inadmissible for conformal p-values as they are outperformed by the Simes method. Indeed, Vovk et al. (2022) demonstrated that universally valid methods are dominated by the Simes method (Simes, 1986), which is applicable to PRDS p-values (Sarkar, 2008). However, beyond scenarios featuring rare and strong signals, even the Simes method is often surpassed by alternative approaches.
As proved in Appendix A1.2, the Simes method applied to conformal p-values is inadmissible because it is dominated by the permutation-based Simes method. However, the difference between the two is minimal unless $m$ and $|S|$ are small, which makes the Simes method appealing for computational reasons.
@@
In general, even if permutation tests are applicable, it is unclear which test statistic should be used in practice. Birnbaum (1954) demonstrated that any combination method is optimal against some alternative. For instance, Fisher’s method, recently refined by Bates et al. (2023) to accommodate the positive dependencies of conformal p-values, is well-suited for scenarios with numerous weak signals (Heard and Rubin-Delanchy, 2018).
@@
Another elegant optimality result is given by Lehmann (1953), which shows that the WMW test is the locally most powerful rank test for the global null hypothesis $\theta=0$ under a non-parametric alternative in which the test scores $Y_{j}$ follow the mixture distribution $(1-\theta)F+\theta F^{2}$. Here, $\theta\in[0,1]$ is the proportion of outliers and $F$ denotes the cumulative distribution function of the calibration scores $X_{i}$. This model assumes the outliers have weak signals, as $F^{2}$ is the distribution of the maximum of two independent random variables with distribution $F$, where $\mathbb{P}(X_{i}<Y_{j})=2/3$ for any specification of $F$. A rare/weak nonparametric model can be obtained for very small values of $\theta$, extending the parametric model that motivates the Higher Criticism test of Donoho and Jin (2004, 2015).
As demonstrated in Appendix A1.4, the Mann-Whitney test statistic for testing $H_{S}$ can be formulated as a combination of conformal p-values: $T^{\text{MW}}_{S}=(m+1)\sum_{j\in S}(1-p_{j})$. Since $T^{\text{MW}}_{S}$ is a monotonic function of $\bar{p}_{S}=|S|^{-1}\sum_{j\in S}p_{j}$, the test statistic can be simplified to the average of conformal p-values $\bar{p}_{S}$, leading to rejection of $H_{S}$ for small values of $\bar{p}_{S}$.
The average of the p-values, also known as Edgington’s method (Edgington, 1972), has been studied by Rüschendorf (1982) and Meng (1994). They show that twice the average of the p-values is a valid p-value for arbitrary dependence. Recently, Choi and Kim (2023) demonstrated that the “twice the average” rule cannot be improved even under the assumption of exchangeability of the p-values.
In contrast, the WMW test, when utilizing the average of the conformal p-values as the test statistic, rejects $H_{S}$ if $\bar{p}_{S}\leq\alpha^{\mathrm{Perm}}$, where the critical value $\alpha^{\mathrm{Perm}}$ is determined by taking the $\alpha$ quantile of the permutation distribution of $\bar{p}_{S}$, ensuring exact validity. This critical value $\alpha^{\mathrm{Perm}}$ may be calculated either exactly in small samples or through asymptotic approximations. A well-known asymptotic approach when $m$ and $|S|$ are both large is that of Hoeffding (1948), which leads to a much more powerful test compared to the combination approach for arbitrary dependence.
@@
In this paper, we focus on applying the ACODE method in conjunction with the Simes test, Fisher’s combination method, and the WMW test, as previewed in Figure 1. These procedures are chosen for their ease of use within our conformal inference framework and for their complementary strengths against a range of alternative hypotheses, including both sparse/strong and dense/weak effects. We detail these methods in Table LABEL:tab:standard_approaches in Appendix A1, where we discuss their underlying assumptions, precision, and the scenarios in which they are most effective. Additional, although non-exhaustive, information about existing local testing procedures can also be found in Appendix A1.
@@
2.3.2 Shirashi’s Locally Most Powerful Rank Test

### v3

@@ Abstract
This paper develops a flexible distribution-free method for collective outlier detection and enumeration, designed for situations in which the presence of outliers can be detected powerfully even though their precise identification may be challenging due to the sparsity, weakness, or elusiveness of their signals. This method builds upon recent developments in conformal inference and integrates classical ideas from other areas, including multiple testing, locally most powerful and adaptive rank tests, and non-parametric large-sample asymptotics. The key innovation lies in developing a principled and effective approach for automatically choosing the most appropriate machine learning classifier and two-sample testing procedure for a given data set. The performance of our method is investigated through extensive empirical demonstrations, including an analysis of the LHCO high-energy particle collision data set.
@@ 1.1 Background and Motivation
While conformal inference is gaining momentum in both statistics and machine learning, its focus in the context of outlier detection has so far primarily been on individual-level outlier identification, where each data point is separately evaluated as a potential outlier. However, individual-level identification is not always feasible. In many applications, outliers may be too rare or weak to reach statistical significance (Donoho and Jin, 2004), or they may exhibit unremarkable behavior in isolation but reveal anomalous patterns when analyzed collectively (Feroze et al., 2021), such as showing under-dispersion relative to inliers.
@@ 1.2 Contributions and Outline
We present ACODE, an Automatic Conformal Outlier Detection and Enumeration method. ACODE leverages two-sample rank tests applied to univariate conformity scores, which can be generated by any model trained to distinguish outliers from inliers. By employing split-conformal inference, ACODE converts these scores into statistical tests controlling the type-I error rate. In addition, by utilizing the closed testing principle (Marcus et al., 1976), ACODE can construct simultaneous lower confidence bounds for the number of outliers within any subset of the test sample (Goeman and Solari, 2011), while also offering a global test for outlier detection as a byproduct of outlier enumeration. While these high-level ideas are intuitive, the flexibility of the conformal inference framework introduces substantial complexity and a great deal of implementation freedom. This brings us to the two key methodological questions addressed in this paper.
@@
Figure 1: Preview of performance of ACODE on the LHCO data. Left: Median and 90th-quantile over repeated experiments for a 90% lower confidence bound for the number of outliers. Right: power against the global null hypothesis of no outliers at the 10% level (horizontal line). The results are shown as a function of the true number of outliers in a test set of cardinality 10,000. ACODE utilizes a testing procedure that may be adaptively selected (red curve) or fixed (other solid curves). Dotted curve: number of individual discoveries or power obtained by applying the Benjamini-Hochberg procedure (BH) to conformal $p$-values, controlling the false discovery rate below 10%.
As detailed in Section 3.2, our analysis of the LHCO data aims to (1) detect outliers, or unusual collision events, within a test sample, and (2) establish a 90% lower confidence bound for the number of these outliers. Our method not only controls errors for both tasks but also provides insightful inferences. This illustrates ACODE’s adaptability, as it selects the most suitable testing procedures in a data-driven manner.
Additionally, Figure 1 highlights the challenge of conducting an adaptive data analysis without incurring selection bias. It compares ACODE’s performance with that of a naive “cherry-picking” approach that greedily tests various classification algorithms and testing procedures, selecting the one with the most appealing outcomes. Unsurprisingly, that can lead to inflated type-I errors because the same data are used both to select the classification algorithm and testing procedure and to assess statistical significance. In contrast, ACODE is protected against selection bias while often achieving similarly high power.
@@
The novel contributions of this paper are the following.
  1. 1.
    We introduce ACODE, a flexible method for outlier enumeration with data-driven selection of the machine-learning classifier and the two-sample testing procedure. This method is built on a novel integration of existing theoretical results on closed testing (Proposition 1), locally most powerful rank tests (Theorem A1), and data-driven adaptivity via sample splitting (Proposition A3), and yields rigorous inferences.
  2. 2.
    We present asymptotic approximations to the null distributions of the locally most powerful rank test and its adaptive version under exchangeable conformity scores (Theorems 1 and A2), extending earlier results that assume i.i.d. scores.
  3. 3.
    We implement an adaptive rank test, which nonparametrically estimates the unknown outlier distribution, that attains the same local asymptotic power as the locally most powerful rank test with oracle knowledge of the outlier distribution (Theorem 2).
@@ 1.3 Related Work
Recent advances in large-scale inference have led to methods for signal detection and estimation of the proportion of non-null effects (Cai and Sun, 2017). In particular, the Higher Criticism test (Donoho and Jin, 2004) has been extended to provide simultaneous bounds on the false discovery proportion (FDP) via closed testing (Goeman et al., 2021). However, the FDP bounds derived from Higher Criticism are valid only under the assumption of independent $p$-values, which does not hold in the conformal inference setting.
@@
In this paper, we integrate conformal inference with closed testing to construct simultaneous lower confidence bounds for the number of outliers in any subset of the test set. A central challenge is selecting an effective local testing procedure suited to the data at hand.
When a uniformly most powerful test is unavailable, a possible strategy is to maximize power locally around the null hypothesis. In a seminal paper, Lehmann (1953) showed that the Wilcoxon–Mann–Whitney (WMW) test (Wilcoxon, 1945; Mann and Whitney, 1947) is the locally most powerful rank test under the alternative in which the outlier distribution is given by the maximum of two independent copies of the inlier distribution. This result was later extended by Shiraishi (1985), who characterized locally most powerful rank tests for arbitrary outlier distributions, but at the cost of requiring their specification.
In practice, however, the assumption that the outlier distribution is known is unrealistic. Rank tests are distribution-free under the null and therefore exact regardless of the underlying distribution, offering maximal protection against model misspecification. Their power, however, depends on the true outlier distribution and may be sensitive to misspecification, potentially yielding a locally optimal test for the wrong model.
This dependence can be mitigated through adaptation. The literature on adaptive rank tests is extensive; here we focus on two basic paradigms, following Sidak et al. (1999). The restrictive paradigm (Hogg et al., 1975) requires to fix a reasonable family of distributions for outliers and select one, based on the data, then perform the corresponding rank test. If the true outlier distribution is close to the defined family, the performed test tends to be nearly optimal. In the non-restrictive paradigm, the outlier distribution is first estimated from the data and then used to construct a test that is optimal for the estimated model. Remarkably, this strategy can succeed: in Theorem 2, we show that the resulting adaptive rank test attains the same local asymptotic power as the locally most powerful rank test that has oracle knowledge of the outlier distribution.
ACODE combines these two approaches by considering both a prespecified family of rank tests and an adaptive rank test based on density estimation, and automatically selecting the procedure with the best empirical performance.
@@
A typical concern with closed testing is the computational cost, generally exponential. However, efficient polynomial-time shortcuts are often available (Goeman et al., 2019; Tian et al., 2023). In Section 2.4 and in Appendix A3 we, respectively, present the closed testing method and the computational shortcuts that make our method feasible and scalable. As we will see, different local tests often require different shortcuts and have different costs.
@@ 2.1 Setup and Problem Statement
$Z_{1},\ldots,Z_{n}\overset{\text{i.i.d.}}{\sim}P_{0},$ $Z_{m+j}\overset{\text{ind.}}{\sim}P_{j},\;\forall j\in[m].$ (1)
In the following, we will refer to the data points sampled from $P_{0}$ as inliers and those from any $P_{j}\neq P_{0}$ as outliers. For each $j\in[m]$, define the null hypothesis $H_{j}:P_{j}=P_{0}$. Let $I_{0}:=\{j\in[n]:P_{j}=P_{0}\}$ denote the subset of inliers in the test set. Similarly, let $I_{1}=[n]\setminus I_{0}$ denote the subset of outliers in the test set.
Prior work on conformal inference for outlier detection (Bates et al., 2023; Marandon et al., 2024) has primarily focused on testing individual-level hypotheses $H_{j}$ for each $j\in[m]$, with the goal of controlling the FDR. However, as discussed in Section 1, there are many scenarios where the power to reject these individual-level hypotheses $H_{j}$ is very low, highlighting the need for an alternative approach.
In this paper, we address two key tasks related to collective outlier detection and enumeration. The first task, outlier detection, involves testing the intersection hypothesis $H_{S}:=\cap_{j\in S}H_{j}$, which posits that a subset $S\subseteq[n]$ of test points contains no outliers. The second task, outlier enumeration, focuses on estimating the number of outliers within a subset $S\subseteq[n]$, specifically by providing a lower confidence bound, without necessarily identifying them. For both tasks, we consider cases where $S$ is either fixed or data-driven. In the latter scenario, we will utilize the closed testing principle to obtain simultaneous inferences that remain valid for all possible subsets $S$ (Goeman and Solari, 2011). This approach gives practitioners flexible tools for interactive data analysis.
@@ 2.2 Computing Conformity Scores and Conformal $p$-values
As long as $(X_{1},\ldots,X_{m},(Y_{i})_{i\in I_{0}})$ are exchangeable, $p_{j}$ is a valid $p$-value for $H_{j}$, in the sense that $\mathbb{P}[p_{j}\leq\alpha\;;\;H_{j}]\leq\alpha$ for all $\alpha\in[0,1]$. Further, $p_{j}$ is not too conservative, in the sense that $\mathbb{P}[p_{j}\leq\alpha\;;\;H_{j}]$ is not far from $\alpha$, as long as $m$ is large and the conformity scores are almost surely distinct; the latter would be a mild assumption (since one can always add a small amount of independent noise to $\hat{s}$) but is not required.
@@ 2.3.1 Existing Approaches
Gazin et al. (2024) characterize the dependence structure of conformal $p$-values, which depends only on the ranks of the conformity scores. Under the null hypothesis, these ranks are exchangeable. Ignoring this structure by using aggregation methods that are valid under arbitrary dependence leads to inadmissible procedures—this follows from Vovk et al. (2022), which show that universally valid methods are dominated by the Simes procedure (Simes, 1986), which is applicable to conformal $p$-values (Sarkar, 2008).
@@
While the Simes method tends to work well in scenarios featuring rare and strong signals, it is itself also inadmissible (see Appendix A1.2), and in practice it is often less powerful than alternative approaches because any $p$-value combination method is optimal against some alternative (Birnbaum, 1954). For instance, Fisher’s method, recently refined by Bates et al. (2023) to accommodate the positive dependencies of conformal $p$-values, is well-suited for scenarios with numerous weak signals (Heard and Rubin-Delanchy, 2018).
@@
A related optimality result is due to Lehmann (1953), which establishes that the WMW test is the locally most powerful rank (LMPR) test for the global null hypothesis $\theta=0$ under a nonparametric alternative in which the test scores $Y_{j}$ follow the mixture distribution $(1-\theta)F+\theta F^{2}$. Here, $\theta\in[0,1]$ is the proportion of outliers and $F$ denotes the cumulative distribution function of the calibration scores $X_{i}$. In this model the outliers have weak signals, as $F^{2}$ is the distribution of the maximum of two independent random variables with distribution $F$, where $\mathbb{P}(X_{i}<Y_{j})=2/3$ for any $F$. As shown in Appendix A1.4, the Mann-Whitney statistic for testing $H_{S}$ can be formulated as a combination of conformal $p$-values: $T^{\text{MW}}_{S}=(m+1)\sum_{j\in S}(1-p_{j})$. Since $T^{\text{MW}}_{S}$ is a monotonic function of $\bar{p}_{S}=|S|^{-1}\sum_{j\in S}p_{j}$, the test statistic can be simplified to the average of conformal $p$-values $\bar{p}_{S}$.
Finally, permutation tests offer significant flexibility, but in addition to their higher computational cost, which makes them less well-suited to be integrated within a closed-testing framework, they suffer from the same limitation as the aforementioned approaches that in practice it is unclear which test statistic should be used for a given data set.
@@
In this paper, we focus on applying ACODE in conjunction with Simes test, Fisher’s combination method, and WMW test, as previewed in Figure 1. These procedures are chosen for their ease of use within our conformal inference framework and for their complementary strengths against a range of alternative hypotheses, including both sparse/strong and dense/weak effects. We detail these methods in Table A1 in Appendix A1, discussing their underlying assumptions, precision, and the scenarios in which they are most effective. Additional, although non-exhaustive, information about existing local testing procedures can be found in Appendix A1.
@@
2.3.2 Shirashi’s Locally Most Powerful Rank (LMPR) Test
In this section, let us imagine the calibration scores $X_{1},\ldots,X_{m}$ are drawn independently from an unknown distribution $F$ and the test scores $Y_{1},\ldots,Y_{n}$ are drawn independently from a mixture distribution $(1-\theta)F+\theta G(F)$, where $\theta\in[0,1]$ is the unknown proportion of outliers and $G$ is a distribution function on $[0,1]$ with density function $g$. The global null hypothesis $H_{[n]}:\cap_{j\in[n]}H_{j}$ of no outliers implies $H_{0}:\theta=0$ under this mixture model.